OpenStudy (anonymous):

Problem : Use the chain rule method and plss show ur solution so i can understand it y=(x^2+5)^5/(x^4+3)^3

OpenStudy (anonymous):

WAT DO U WANT

OpenStudy (calculusfunctions):

You would like to differentiate y w.r.t. x, correct?

OpenStudy (calculusfunctions):

You can use the chain rule within the quotient rule.

OpenStudy (anonymous):

\[y'=(x ^{2}+5)^{5}\div(x ^{4}+3)^{3}\]

OpenStudy (anonymous):

im confuse.

OpenStudy (calculusfunctions):

The given function is the derivative function? So the first time, you typed it incorrectly? There seem to be two options here. Either find the derivative or the antiderivative (integral)? Which is it? No one can help you unless you make your intentions clear by typing the question properly.

OpenStudy (anonymous):

derivative sir

OpenStudy (calculusfunctions):

You mean you're given\[\frac{ dy }{ dx }=\frac{ (x ^{2}+5)^{5} }{ (x ^{4}+3)^{3} }\]and you would like to find the anti-derivative. Is that correct?

OpenStudy (anonymous):

the derivative

OpenStudy (calculusfunctions):

So you mean it's\[y =\frac{ (x ^{2}+5)^{5} }{ (x ^{4}+3)^{3} }\]and you would like to find the derivative? Which is it?

OpenStudy (anonymous):

using chain rule . find derivative y'

OpenStudy (calculusfunctions):

OK then finally we're getting somewhere! Do you know the quotient rule? Quotient Rule: If y = f/g then\[y \prime =\frac{ f \prime g -fg \prime }{ g ^{2} }\]Do you know this rule?

OpenStudy (anonymous):

i know the method but im really confused

OpenStudy (calculusfunctions):

About what?

OpenStudy (anonymous):

all this problem

OpenStudy (calculusfunctions):

OK we have\[y =\frac{ (x ^{2}+5)^{5} }{ (x ^{4}+3)^{3} }\]Let's take baby steps and start by letting\[f =(x ^{2}+5)^{5}\]and\[g =(x ^{4}+3)^{3}\]Now do you know how to find f' and g'?

OpenStudy (anonymous):

\[f=4(x ^{2}+5)^{4}(2x)\]

OpenStudy (calculusfunctions):

No. Try again!

OpenStudy (anonymous):

\[f=5(x ^{2}+5)^{4}(2x)\]

OpenStudy (calculusfunctions):

Yes you can!! See! Very good except that it's f'. You wrote f.

OpenStudy (calculusfunctions):

Now find g' similarly.

OpenStudy (anonymous):

is (2x) have exponent of ^5 ??

OpenStudy (calculusfunctions):

No.

OpenStudy (anonymous):

\[g=3(x ^{4}+3)^{2}(4x ^{3})\]

OpenStudy (anonymous):

ahh ok

OpenStudy (calculusfunctions):

Yes and you mean g'. Why do you keep refusing to indicate the prime for derivative? Without that one simple symbol, it's wrong. Understand?

OpenStudy (anonymous):

ah sorry im just lazy typing that :)

OpenStudy (calculusfunctions):

OK now substitute f, f', g, and g' into the quotient rule formula I posted above.

OpenStudy (calculusfunctions):

Well, as a teacher I'm telling you it's wrong without that symbol and as a teacher it's really irritating to see you do that time and time again, even after I tell you not to. I would give you a zero just for that.

OpenStudy (anonymous):

\[y'=\frac{ 5(x ^{2}+5)^{4}(2x) }{ 3(x ^{4}+3)^{2}(4x ^{3}) }\]

OpenStudy (calculusfunctions):

NO! That's not the formula I gave you! Stop being so lazy and look!!

OpenStudy (anonymous):

ahh i know it

OpenStudy (anonymous):

pellet im still confused

OpenStudy (calculusfunctions):

\[f=(x ^{2}+5)^{5}\]\[g =(x ^{4}+3)^{3}\]\[f \prime =5(x ^{2}+5)^{4}(2x)\]\[g \prime =3(x ^{4}+3)^{2}(4x ^{3})\]Now\[y \prime =\frac{ f \prime g -fg \prime }{ g ^{2} }\]You're telling me you're confused about plugging things into a formula?? Really??

OpenStudy (anonymous):

ok i should just solve this using this formula

OpenStudy (calculusfunctions):

Not solve, SIMPLIFY!!

OpenStudy (anonymous):

ok just calm down .. huhuhu in just a noob

OpenStudy (anonymous):

*im

OpenStudy (calculusfunctions):

Don't expand the denominator though, only simply the numerator, preferably by factoring.

OpenStudy (calculusfunctions):

If properly expressed, the numerator should be in factored form in the final step. Good luck! I have to go now.

OpenStudy (anonymous):

tnx